Mickley Company’s plantwide predetermined overhead rate is $22.00 per direct labor-hour and its direct labor wage rate is $15.00 per hour. The following information pertains to Job A-500: Direct materials $ 300 Direct labor $ 150 Required: 1. What is the total manufacturing cost assigned to Job A-500? 2. If Job A-500 consists of 90 units, what is the unit product cost for this job? (Round your answer to 2 decimal places.)
Answer #1

Total Direct labor hours for job A - 500 = Direct labor cost / direct labor wage rate
                                                                   = $150/ $15 per hour
                                                                   = 10 hrs

Total over head cost = overhead cost per labor hours * no. of labor hours
                                 = $22 * 10 hrs
                                 = $220

1) Total manufacturing cost of Job A-500 = Direct materials cost + Direct labor cost + Total over head cost
                                                  = $300 + $150 + $220
                                                  = $670

2) unit product cost for Job A-500 (for 90 units ) = total manufacturing cost/ number of units
                                                                                 = $670/90 units
                                                                                 = $7.44 per unit (round to two Decimal)